40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6138     Entry time: Mon Dec 19 23:50:23 2011
Author: Zach 
Type: Update 
Category: RF System 
Subject: RAMmon 

I have been looking at the swings in the RAMmon channels since the heater was reengaged, to compare them to the data from beforehand (with and without the foam box). With the large grains of salt that I will list after, it appears that the EOM temperature controller does in fact reduce the amplitude of the swings by a measurable factor.


  1. The reason I have not included any plots here is because the data suck. What we should ideally have is a continuous stream of RAMmon signals split into three chunks: 1) no foam, no heat, 2) foam, no heat, and 3) foam and heat. Instead, we have pieces of each kind of data on different days, before and after the MC has been realigned, some in old channels and some in new so that the calibration is different, etc. This piecemeal shit will not do.
  2. I realized that the LF boost was not engaged on the heater when I turned it on most recently. For this reason, the EOM temperature has not been stabilized as well as it might have been on diurnal timescales, and so with the boost it could be that the noise reduction is greater. For posterity, the DC suppression level is ~20x without the boost.

It seems impractical to try and rope off essentially 3 straight days where nothign major can be done to the IFO just to take RAM data. Instead, I think we should figure out a way to mimic the diurnal temperature swings on ~hour timescales. The EOM can temperature follows PSL-FSS_RMTEMP almost exactly and with a very short delay, so we can probably even accomplish this by stepping the lab A/C temperature. If this won't work, we can use an incandescent lamp or something similar to heat up the area around the EOM by a noticeable amount.

I'll try to come up with a good way to do this so that we can get some reliable data...

ELOG V3.1.3-